A right-circular cone is inscribed in a sphere having a fixed radius of 10 in. Express the volume of the cone as a function of its radius.
1
Expert's answer
2020-09-22T12:57:03-0400

The volume of the right circular cone is calculated by the formula

"V = \frac 1 3 \pi r^{2}H"

where r is the radius and H height right-circular cone and R=10 sphere radius

"H = R+\sqrt{R^{2}-r^{2}}"

or

"H = R-\sqrt{R^{2}-r^{2}}"



hence the function of the volume of the cone from its base

"V = \frac 1 3 \pi r^{2}H = \frac 1 3 \pi r^{2}*( R±\sqrt{R^{2}-r^{2}}) =\frac 1 3 \pi r^{2}*( 10±\sqrt{100-r^{2}})"


Answer : "V= \frac 1 3 \pi r^{2}*( 10±\sqrt{100-r^{2}})"